# High-speed ditching of double curvature specimens with cavitation and ventilation

### v1

@@ Abstract
The water entry at high horizontal speed of double-curvature specimens, reproducing the rear part of the fuselage that first gets in contact with the water during aircraft ditching, is investigated experimentally. Three shapes are analysed, representing different aircraft types. Pressure and load measurements are taken, supported by underwater high-speed visualization. The effects of different horizontal speeds, pitch angles, and curvatures are analysed, while keeping a constant vertical-to-horizontal velocity ratio. It is observed that the longitudinal curvature plays a significant role in the hydrodynamics, potentially leading to cavitation and ventilation in the rear part of the specimen at high speeds. The transverse curvature affects the pressures and loads both at the front and at the rear, since a lower transverse curvature increases the possibility of the fluid to escape from the sides. An increase in pitch angle results in an increased loading, in terms of both amplitude and distribution. Finally, an increased horizontal speed leads to higher loads at the front but also affects the cavitation and ventilation modalities at the rear. The load scaling for flat plates also applies to these specimens at the front, but the occurrence of cavitation and ventilation invalidates it at the rear. The time evolution of the cavitation region and of the wetted area are examined through advanced image processing techniques and pressure data analysis. The comparison with the geometric intersection area of the specimen with the still water surface provides further insight into the hydrodynamic phenomena during the water entry.
@@ 1 Introduction
Aircraft ditching is the manoeuvre performed when the pilots and crew are forced to recur to an emergency water landing due to exceptional circumstances such as engine failure, fuel exhaustion, severe weather or other critical and unexpected situations. The goal of ditching is to land the aircraft safely on the water surface, allowing the occupants to be evacuated and rescued with the lowest risk of injury. Although being a rare event, airworthiness regulations require a precise certification for ditching, including demonstration tests proving that the aircraft structural damage during this eventuality is limited and that a sufficient floating time is guaranteed for a safe evacuation of the occupants (Climent et al., 2006).
The numerical approach for certification is becoming feasible but is is still in a developing phase (Siemann & Langrand, 2017; Siemann et al., 2018; Streckwall et al., 2007; Guo et al., 2013; Hughes et al., 2013; Bisagni & Pigazzini, 2018; Anghileri et al., 2011, 2014; Xiao et al., 2017), mainly because of the high resolution needed to correctly capture the hydrodynamic and the fluid-structure interaction phenomena, which implies quite large computational efforts. In order to ascertain the accuracy of the computational models and the reliability of the results, a careful validation and verification is deemed strongly necessary.
The experimental investigation allows to correctly reproduce the physical phenomena occurring during a ditching event. However, ditching tests of a full aircraft are impractical and also too expensive. Attempts to perform experiments mimicking the ditching scenario have been made during the last decades, starting from the pioneering work of Smiley (1951) who performed tests on flat plates entering water with a relevant horizontal speed. Smiley was able to retrieve accurate measurements of pressures, loads and of the wetted area at different test conditions. Even if remarkable for that time, this work has important limitations especially in terms of the data sampling frequency which, being of order of 1 kHz, does not enable to capture the fast transients occurring in a water entry problem (Van Nuffel et al., 2013). Another notable experimental ditching campaign was performed by McBride & Fisher (1953), who conducted free fall tests of several scaled fuselage models, representative of different types of aircraft. The study mainly focused on the role of the longitudinal and transverse body curvatures on the dynamics of the aircraft during ditching. The main limitation of the study lies in the impossibility in the control the impact conditions. In addition, information on the hydrodynamics developing beneath the models were provided, where it is expected that suction forces and possibly even ventilation phenomena may develop, as shown in Tassin et al. (2013).
Scaled model test of a full aircraft model were performed in Climent et al. (2006) and in Zhang et al. (2012). However, the commonly used scaling based on the Froude approximation does not apply in this case, in which highly non-linear hydrodynamic phenomena take place. Furthermore, the aircraft models used in the tests were too rigid, and therefore the fluid-structure interaction aspect could not be correctly reproduced. Focusing on hydrodynamic aspects only, since viscous effects are known to have a negligible role during water entry (Abrate, 2011; Facci et al., 2015; Moghisi & Squire, 1981), the similarity in terms the ventilation and cavitation numbers (Brennen, 2014) is deemed very critical. Such similarity conditions could be achieved only by performing tests in a de-pressurized environment.
In order to overcome such limitations, a High Speed Ditching Facility (HSDF) was built at the CNR - Institute of Marine Engineering. The HSDF allows tests at horizontal speed up to 48 m/s, vertical to horizontal velocity ratios ranging from 0.03 to 0.05 and on test models about 0.7 m wide and 1.3 m long. The pitch angle of the impacting body can be varied from 4 to 10. Depending on the needs, pressures, strains and forces acting on the body can be measured through a ruggedized acquisition system, hosted inside an acquisition box firmly connected to body. High-speed underwater videos are also recorded to help the interpretation of the measurements and to provide additional insights about the physical phenomena taking place during the impact (Iafrati et al., 2015).
@@
Early tests were performed on a flat aluminium plates, 15 mm thick, in a configuration similar to that used in Smiley (1951). Different test conditions were examined by varying the horizontal speed, the vertical-to-horizontal velocity ratio and the pitch angle. The strong correlation between the spray root propagation velocity and the intensity of the pressure peak was confirmed and a scaling of the total loading with respect to the impacting velocity relative to the body was observed (Iafrati, 2015, 2016). Significant fluid-structure interaction phenomena were observed by carrying out tests of thinner aluminium plates, 3 mm and 0.8 mm thick. Differently from the expectations, it was found that higher hydrodynamic loads, up to 30%, occur in presence of large deformations of the plates and, therefore, the estimate of the loads based on tests on rigid specimens is not always conservative (Spinosa & Iafrati, 2021).
@@
Although the tests on flat plates provided a wide dataset for the validation of the capabilities of the computational models to describe the hydrodynamics and the strong fluid-structure interaction phenomena occurring during ditching, the lower portion of an aircraft fuselage is characterised by a curved shape. In order to assess the effects of the transverse curvature of the fuselage, tests on both concave and convex plates, still 15 mm thick, were performed (Iafrati, 2018). However, the results of some numerical simulations and experimental tests on full fuselages showed that the longitudinal curvature at the rear may generate suction forces, associated with negative pressures, which affect the aircraft dynamics during the ditching (Climent et al., 2006; Zhang et al., 2012). In order to better investigate the phenomenon, an experimental test campaign on double curvature rigid specimens was conducted. From a preliminary analysis of the data it was found that cavitation and ventilation may occur at high speed and they may have important effects on the magnitude and distribution of the loading (Iafrati & Grizzi, 2019; Iafrati et al., 2020).
In the present work, a deeper analysis of the tests performed on double curvature specimens is provided. The objective of the analysis is to investigate the role played by the various parameters of the tests, specifically horizontal velocity, pitch angle and body curvatures. The analysis is presented in terms of pressures and total loads acting on the impacting specimens. The availability of the data of the loads at both rear and front makes it possible to identify the significant changes in the load distribution, when cavitation and ventilation phenomena occur. The evolution in time of the wetted area and of the cavitation zone is also investigated based on underwater visualisations, and quantitative estimates are also provided.
@@ 2.1 The High-Speed Ditching Facility
The experiments are performed in the High Speed Ditching Facility (HSDF) of the CNR - Institute of Marine Engineering, described in detail in Iafrati et al. (2015), in which the test repeatability is also assessed. For the purpose of the present paper, it is worth recalling the most relevant features.
@@ 2.2 Specimen shapes
The coordinates associated with the ground-fixed reference frame are indicated with upper-case letters $X$, $Y$ and $Z$. In particular, $X$ is parallel to the undisturbed free surface, $Z$ is vertical and directed upwards and $Y$ is in the direction on a right-handed coordinate system. The specimen horizontal and vertical speed are defined in this reference frame as $U=dX/dt$ and $V=-dZ/dt$, respectively. The coordinates associated with the trolley-fixed reference frame are indicated with lower-case letters $x$, $y$ and $z$, with origin O. As shown in Figure 4 the origin is located at the leading edge of the plate. The $x$ coordinate is parallel to the longitudinal axis of the trolley, the $z$-coordinate is normal to it and directed upwards, whereas the $y$ axis forms and right-handed coordinate system. The pitch angle $\alpha$ is defined as the angle between the trolley axis $x$ and the $X$-direction, whereas the angle $\beta=\arctan{V/U}$ is introduced to denote the angle between the the impact trajectory of the specimen and the undisturbed water level.
@@ 2.3 On-board Measurements
Pressures are acquired through thirty pressure probes type Kulite XTL 123B, full scale range of 300 psi, properly calibrated before starting the test campaign. The probes are installed flush to the external surface of the specimen with $\pm$ 0.1 mm accuracy. The position of the probes is given in Figure 5. In the analysis of the data, beside the specific probe number, reference is also made to probe lines ($l1,l2,l3,l4$) and rows ($r1,r2,...,r8$) as indicated in the figures.
@@
Figure 5: Position of the pressure probes and definition of the row and lines used for the discussion. Dimensions are in mm.
The pressure time histories presented in the following are raw data, unless otherwise specified. It is worth noting that sometimes pressure spikes of non-physical origin may appear, which, among other reasons, are due either to slightly protruding probes generating at such speeds cavitating vortices (Iafrati et al., 2015) or to the temperature shock experienced by the sensors when they get in touch with water (Van Nuffel et al., 2013), which is at a lower temperature than air.
@@
Figure 6: Position of the load cells, referring to the specimen S2. The total rear and front force direction is indicated, as well as the trolley-fixed reference frame. The dimensions are in mm. The value of $H$ is 520 mm, 440 mm and 436.2 mm for the shapes S1B, S2 and S3, respectively.
Four Kistler 9343 cells, with full scale range of 70 kN, are used in the $z$-direction, whereas two Kistler 9363 cells, with full scale range of 120 kN, are used in the $x$-direction. The load cell signals are conditioned using a charge amplifier type Kistler 5073-A421. The channels are connected in a way that the loads acting in the $z$-direction can be distinguished in the rear $F_{zR}$ and in the front $F_{zF}$ contributions, whereas $F_{x}$ denotes the load acting in the $x$-direction. The distance $H$ for the three specimens S1B, S2 and S3 are 520 mm, 440 mm and 436.2 mm respectively. Although the single load cell is very accurate when loaded directly along its axis, the connection to the frame of the acquisition box introduces some uncertainties, since the loading is not exactly aligned with the axes of the cells. Furthermore, due to the inherent free play of the installations, some errors may be expected for low values of the loading. For this reason a careful calibration was carried out before starting the test campaign, and it was found that the recorded data are always below the actual loading. The underestimate ranges from about 8% for a load of order of 8 kN and reduces continuously when increasing the load, being about 4% for a load of order of 14 kN. The load was also applied at different locations along the $x$-axis and the centre of loads could be retrieved with a satisfactory accuracy by the combined data of the rear and the front load cells. Unless specified differently, also the force time histories provided in the following are raw data.

### v2

@@ Abstract
The water entry at high horizontal speed of double-curvature specimens, representing the rear part of the fuselage that first contacts the water during aircraft ditching, is investigated experimentally. Three shapes, reproducing different aircraft types, are analysed through pressure and load measurements, supported by underwater high-speed visualizations. Tests are performed at different horizontal speeds and pitch angles, while maintaining a constant vertical-to-horizontal velocity ratio. The longitudinal curvature is found to significantly affect the hydrodynamics, potentially leading to cavitation and ventilation at the rear at high speeds. The transverse curvature influences pressures and loads at both the front and rear, with higher transverse curvature favouring lateral fluid escape. Higher pitch angles increase loading in both magnitude and distribution. Increasing the horizontal speed results in higher loads at the front and alters cavitation and ventilation modalities at the rear. Load scaling laws established for flat plates are valid at the front but are invalidated at the rear due to cavitation and ventilation. The time evolution of the cavitation and wetted area is analysed through advanced image processing and pressure data. Comparison with the geometric intersection area between the specimen and the still water level offers further insight into the water entry hydrodynamics.
@@ 1 Introduction
Aircraft ditching is the manoeuvre performed when the pilots and crew are forced to recur to an emergency water landing due to exceptional circumstances such as engine failure, fuel exhaustion, severe weather or other critical unexpected situations. The goal of ditching is to land the aircraft safely on the water surface, allowing the occupants to be evacuated and rescued with the lowest risk of injury. Although being a rare event, airworthiness regulations require a precise certification for ditching, including demonstration tests proving that the aircraft structural damage during this eventuality is limited and that a sufficient floating time is guaranteed for a safe evacuation of the occupants, as specified in European Union Aviation Safety Agency (2020) and Federal Aviation Administration (2023) (see also Desjardins et al. (1989)).
The theoretical foundations of ditching were laid by Von Karman (1929) and Wagner (1932), which first established a theoretical framework of vertical water entry. These initial (semi)-analytical models were successively refined by several researchers (Logvinovich, 1969; Scolan and Korobkin, 2001; Korobkin and Scolan, 2006; Korobkin, 2004; Tassin et al., 2010). forming the basis of low-fidelity numerical approaches, which became popular in the early computational era. A notable example is the Boundary Element Method (BEM) (Zhao and Faltinsen, 1993; Battistin and Iafrati, 2003; Iafrati and Korobkin, 2004, 2008), based on potential theory. Owing to their simplified discretization and rapid computation, these methods are still used today for quick, approximate solutions.
To extend vertical entry solutions to the ditching case, where a significant horizontal velocity is also present, the multi-section or the 2D+t approach can be used (Fontaine and Tulin, 1998), which break down the 3D motion into a series of simpler 2D cross-section analyses (Sun and Faltinsen, 2011), albeit with some limitations. The application of a multi-section approach to aircraft ditching is presented in Del Buono et al. (2021); Del Buono and Iafrati (2025). The DITCH code (Bensch et al., 2003; Streckwall et al., 2007) implements a 2D+t method, complemented by empirical formulations and has also been recently used with machine learning to predict ditching loads (Schwarz et al., 2025).
As computing power increased, high-fidelity methods using finer discretization of the Navier–Stokes equations were increasingly employed. Early high-fidelity tools are reviewed in Seddon and Moatamedi (2006) and Hughes et al. (2013). These methods often combine fluid dynamics solvers with structural models, like Finite Element Methods (FEM), for fluid–structure interaction. A common approach uses URANS solvers with Volume-Of-Fluid (VOF) to capture the free surface evolution. This method is applied to aircraft ditching in (Guo et al., 2013; Qu et al., 2015, 2016; Wang, 2023; Zheng et al., 2021, 2025). Free surface evolution can also be captured using over-set approaches (Zha et al., 2024; Song et al., 2023) and Spinosa et al. (2022), the latter focusing on the final phase of ditching after water impact. The ALE method overcomes mesh deformation in Lagrangian methods via periodic re-meshing (Cheng et al., 2011; Siemann and Langrand, 2017; Bisagni and Pigazzini, 2018; Feng et al., 2020; Fan et al., 2024). The CEL method uses a fixed Eulerian mesh for fluids and a moving Lagrangian mesh for solids, to model fluid–structure interaction (Goron et al., 2023, 2025). Mesh-free methods like Smoothed Particle Hydrodynamics (SPH) avoid mesh distortion and capture sprays well. However, they require higher computational resources and still face challenges in modelling cavitation during ditching (Anghileri et al., 2011, 2014; Groenenboom and Cartwright, 2010; Groenenboom and Siemann, 2015; Siemann and Langrand, 2017; Siemann et al., 2018; Bisagni and Pigazzini, 2018; Groenenboom et al., 2021; Muñoz et al., 2025; Wang et al., 2024). Undoubtedly, numerical simulations of ditching have advanced significantly. However, their use in ditching certification remains limited due to high computational costs and the need for rigorous validation against experimental data.
Experimental investigations enable an accurate replication of the physical phenomena in a ditching event and are essential for validating the various computational tools. However, full-scale tests of an entire aircraft remain impractical and prohibitively expensive. Attempts to replicate ditching through scaled-model experiments began in the 1950s with pioneering work by Smiley (1951), who tested flat plates entering water at significant horizontal speeds, obtaining accurate measurements of pressures, loads, and wetted areas under various conditions. While remarkable for its time, the study had limitations, particularly a data sampling frequency of around 1 kHz, insufficient to capture the rapid transients typical of water entry problems. Another notable experimental ditching campaign from the same period was conducted by McBride and Fisher (1953), who performed free-fall tests on various scaled fuselage models representing different aircraft types. The study primarily examined the effects of longitudinal and transverse body curvatures on ditching dynamics. Its main limitation was the inability to control impact conditions. A review of early experimental ditching activities is provided by Smith et al. (1957), while more recent developments are covered in Seddon and Moatamedi (2006); Abrate (2011). More recent scaled model tests of a full aircraft model were presented in Climent et al. (2006); Zhang et al. (2012); Song et al. (2023); Wenli et al. (2025).
The process of scaling aircraft ditching tests from full to model scale presents substantial challenges. In ship hydrodynamics, towing tank experiments often rely on Froude similarity, which ensures correct scaling of the ratio between inertial and gravitational forces. When Froude scaling is applied, the model test speed decreases proportionally to the square root of the scale factor. In the hydrodynamics of water entry, the Reynolds number, representing the ratio of inertial to viscous forces, should also be considered, however, previous studies have shown that viscous effects are negligible (Facci et al., 2015; Moghisi and Squire, 1981). Aircraft ditching involves high horizontal velocities and large-scale impact phenomena, which generate extremely high pressures in some regions of the wetted surface and very low pressures in others. Under these conditions, suction, cavitation and ventilation are likely to occur Hughes et al. (2013); Climent et al. (2006); Zhang et al. (2012); Tassin et al. (2013) Therefore, achieving similarity in terms of cavitation and ventilation numbers is essential (Brennen, 2014). This type of similarity can only be reproduced through testing in a de-pressurized environment or by reaching sufficiently high speeds.
To address these challenges, the High-Speed Ditching Facility (HSDF) at the CNR- Institute of Marine Engineering was designed and built. The HSDF allows tests at horizontal speeds of up to 48 m/s, with vertical-to-horizontal velocity ratios between 0.03 and 0.05, using models approximately 0.7 m wide and 1.3 m long. This setup enables near full-scale hydrodynamic conditions and allows observation of phenomena such as cavitation and ventilation, which cannot be captured at lower speeds or smaller scales. Pressures and loads can be acquired at high sampling rates through an integrated on-board measurement system, while high-speed underwater imaging provides complementary insight into the flow phenomena (Iafrati et al., 2015).
@@
Early tests were performed on a flat aluminium plates, 15 mm thick, in a configuration similar to that used in Smiley (1951). Different test conditions were examined by varying the horizontal speed, the vertical-to-horizontal velocity ratio and the pitch angle. The strong correlation between the spray root propagation velocity and the intensity of the pressure peak was confirmed and a scaling of the total loading with respect to the impacting velocity relative to the body was observed (Iafrati, 2015, 2016). The high speeds achievable also allow for a more accurate investigation of fluid–structure interaction on thin aluminium plates (3 mm and 0.8 mm thick). Experimental results showed that hydrodynamic loads can increase by up to 30% in the presence of significant structural deformation. As a result, load estimates based on tests with rigid specimens may not always be conservative (Spinosa and Iafrati, 2021).
@@
Flat plate tests provided validation data for ditching models. However, the curved geometry of the fuselage lower section required further investigation. Transverse curvature effects were examined using 15 mm thick concave and convex plates (Iafrati, 2018). Furthermore, the longitudinal curvature of the fuselage rear portion induces suction forces due to negative pressures, significantly influencing the aircraft ditching dynamics, as demonstrated by numerical and experimental studies (Climent et al., 2006; Zhang et al., 2012). To elucidate this phenomenon, an experimental campaign was conducted on rigid specimens exhibiting double curvature. From a preliminary analysis of the data it was found that cavitation and ventilation may occur at high speed and they may have important effects on the magnitude and distribution of the loading (Iafrati and Grizzi, 2019; Iafrati et al., 2020).
This paper presents additional results from this extensive experimental campaign. Continuing our past research, we further explore the influence of horizontal speed and offer considerations on cavitation modalities and cavitation number. A key objective of this paper is also to investigate the effects of pitch angle and the longitudinal and transverse curvatures, aspects not addressed in previous studies. Analysis of pressures and total loads on the impacting specimens reveals significant effects at the front and rear, linked to cavitation and ventilation phenomena. Additionally, the temporal evolution of the wetted area and cavitation zone is examined through underwater visualizations, with quantitative estimates provided. The complexity of the underlying physics often results in theoretical and numerical models exhibiting limited accuracy, especially at high speeds. This paper offers a physical interpretation alongside a comprehensive database to support the validation of these models.
@@ 2.1 The High-Speed Ditching Facility
The experiments are performed in the High Speed Ditching Facility (HSDF) of the CNR - Institute of Marine Engineering, described in detail in Iafrati et al. (2015). For the purpose of the present paper, it is worth recalling the most relevant features.
@@ 2.2 Specimen shapes
The coordinates associated with the ground-fixed reference frame are indicated with upper-case letters $X$, $Y$ and $Z$. In particular, $X$ is parallel to the undisturbed free surface, $Z$ is vertical and directed upwards and $Y$ is in the direction on a right-handed coordinate system. The specimen horizontal and vertical speed are defined in this reference frame as $U=dX/dt$ and V=dZ/dtV=-dZ/dt, respectively. The coordinates associated with the trolley-fixed reference frame are indicated with lower-case letters $x$, $y$ and $z$, with origin O. As shown in Figure 4 the origin is located at the leading edge of the plate. The $x$ coordinate is parallel to the longitudinal axis of the trolley, the $z$-coordinate is normal to it and directed upwards, whereas the $y$ axis forms and right-handed coordinate system. The pitch angle $\alpha$ is defined as the angle between the trolley axis $x$ and the $X$-direction, whereas the angle $\beta=\arctan{V/U}$ is introduced to denote the angle between the the impact trajectory of the specimen and the undisturbed water level.
@@ 2.3 On-board Measurements
Pressures are acquired through thirty pressure probes type Kulite XTL 123B, full scale range of 300 psi. The probes were carefully calibrated by the manufacturer prior to delivery and were subsequently re-calibrated for verification in the metrology laboratory of CNR-INM before installation. It was found that the sensitivity values were, in all cases, very close to those provided by the manufacturer, with a discrepancy of approximately 1%. Following standard procedures, the relative expanded uncertainty attributed to random effects was determined to be 1.7% of the average measured pressure at a 95% confidence level and 0.9% at a 67% confidence level. The probe was also tested at a few temperatures within its 0–40 °C compensation range, showing a linear sensitivity of about 0.021 kPa/°C. A typical 5 °C temperature change, experienced as the probes operate between approximately 20 °C and 15 °C, causes a negligible pressure shift of 0.105 kPa (0.0051% of full scale). This drift is expected and does not require any further compensation. The probes are installed flush to the external surface of the specimen with $\pm$ 0.1 mm accuracy. The position of the probes is given in Figure 5. In the analysis of the data, beside the specific probe number, reference is also made to probe lines ($l1,l2,l3,l4$) and rows (r1,r2,,r8r1,r2,...,r8) as indicated in the figures.
@@
Figure 5: Position of the pressure probes and definition of the row and lines used for the discussion. Dimensions are in $m$.
The pressure time histories presented in the following are raw data, unless otherwise specified. It is worth noting that sometimes pressure spikes of non-physical origin may appear, caused, among other reasons, either by slightly protruding probes generating cavitating vortices at such speeds (Iafrati et al., 2015) or by the temperature shock experienced by the sensors when they are suddenly immersed in water at a lower temperature than air (Van Nuffel et al., 2013). It is important to clarify that the presence of these spikes is not in contradiction with the considerations on temperature sensitivity made above, since they result from sudden temperature shocks, whereas sensitivity tests are performed by manually applying temperature variations, which are consequently much more gradual.
@@
Figure 6: Position of the load cells, referring to the specimen S2. The total rear and front force direction is indicated, as well as the trolley-fixed reference frame. The dimensions are in $m$. The value of $H$ is 520 mm, 440 mm and 436.2 mm for the shapes S1B, S2 and S3, respectively.
Four Kistler 9343 cells, with full scale range of 70 kN, are used in the $z$-direction, whereas two Kistler 9363 cells, with full scale range of 120 kN, are used in the $x$-direction. The load cell signals are conditioned using a charge amplifier type Kistler 5073-A421. The channels are connected so that the loads acting in the $z$-direction can be separated into the rear $F_{zR}$ and front $F_{zF}$ contributions, whereas $F_{x}$ denotes the load acting in the $x$-direction. The distance $H$ for the three specimens S1B, S2 and S3 are 520 mm, 440 mm and 436.2 mm respectively. Although the single load cell is very accurate when loaded directly along its axis, the connection to the frame of the acquisition box introduces some uncertainties, since the loading is not exactly aligned with the axes of the cells. Furthermore, due to the inherent free play of the installations, some errors may be expected for low values of the loading. For this reason a careful calibration was carried out before starting the test campaign, and it was found that the recorded data are always below the actual loading. The underestimate ranges from about 8% for a load of order of 1.2 kN and reduces continuously when increasing the load, being about 4% for a load of order of 14 kN. The load was also applied at different locations along the $x$-axis and the centre of loads could be retrieved with a satisfactory accuracy by the combined data of the rear and the front load cells. Unless specified differently, also the force time histories provided in the following are raw data.
@@
The repeatability of the pressure and load measurements was thoroughly investigated by Iafrati et al. (2015) in the context of flat plate tests conducted in the HSDF. These considerations are directly applicable to the present experiments, as the same types of pressure probes, load cells and data acquisition systems were employed. Nevertheless, to verify the experimental repeatability under the current conditions, each test was repeated at least twice. The resulting time histories showed excellent agreement, with only minor discrepancies. These were primarily attributed to the inherent difficulty in precisely reproducing the same target nominal velocity, which led to a small variation of approximately 1%.